tan60°-tan30°---------------------=tan30°1+tan60°. tan30°

Grade:10

4 Answers

CHAKRAVARTHI
6 years ago
Answer is” tan-1[1+ square root of 3] /2
Given ,
 = tan600-[tan300][X]=tan300+tan600.tan300
=square root of 3-1/ square root of 3[X]=1/ square root of 3+[ square root of 3.1/ square root of 3]  …. [CANCEL]
=[3-1] / square root of 3[X]=1/ square root of 3 +1
=finally we get,
 X =[1+ square root of 3] /2
CHAKRAVARTHI
6 years ago
 
Answer is” tan-1[1+ square root of 3] /2
Given ,
 = tan600-[tan300][X]=tan300+tan600.tan300
=square root of 3-1/ square root of 3[X]=1/ square root of 3+[ square root of 3.1/ square root of 3]  …. [CANCEL]
=[3-1] / square root of 3[X]=1/ square root of 3 +1
=finally we get,
 X =[1+ square root of 3] /2
CHAKRAVARTHI
6 years ago
 
Answer is” tan-1[1+ square root of 3] /2
Given ,
 = tan600-[tan300][X]=tan300+tan600.tan300
=square root of 3-1/ square root of 3[X]=1/ square root of 3+[ square root of 3.1/ square root of 3]  …. [CANCEL]
=[3-1] / square root of 3[X]=1/ square root of 3 +1
=finally we get,
 X =[1+ square root of 3] /2
 
CHAKRAVARTHI
6 years ago
 
Answer is” tan-1[1+ square root of 3] /2
Given ,
 = tan600-[tan300][X]=tan300+tan600.tan300
=square root of 3-1/ square root of 3[X]=1/ square root of 3+[ square root of 3.1/ square root of 3]  …. [CANCEL]
=[3-1] / square root of 3[X]=1/ square root of 3 +1
=finally we get,
 X =[1+ square root of 3] /2